OpenStudy (anonymous):

5/8 of a persons money is 367.50 what is the total amount of money the person has

OpenStudy (anonymous):

If 5/8=367.50 Then 1/8 is (367.50)/5 1/8=73.50 To get the whole figure, (1/8) * 8 gives you 73.50*8 Total amount=588.00

OpenStudy (anonymous):

you are a live saver and a genius..i really appreciate it..cheers man..